Completely factor this polynomial:
-------------------------
4x² - 25
---------------------------
i need to know the steps on how to solve it

Found 3 solutions by Vladdroid, stanbon, lynnlo:
Answer by Vladdroid(91)

FACTORED
4x^2-25
a^2-b^2=(a-b)(a+b)
(2x-5)(2x+5)==========answer

Answer by stanbon(75887)

4x² - 25
=========
= (2x-5)(2x+5)
------------------
Cheers,
Stan H.

Answer by lynnlo(4176)

FACTORED
4x^2-25
a^2-b^2=(a-b)(a+b)
(2x-5)(2x+5)==========answer